# TaoSR-AGRL: Adaptive Guided Reinforcement Learning Framework for E-commerce Search Relevance

### v2

@@ Abstract.
Query–product relevance prediction is fundamental to e-commerce search and has become even more critical in the era of AI-powered shopping, where semantic understanding and complex reasoning directly shape the user experience and exert an indirect, yet substantial impact on business conversion. Large Language Models (LLMs) enable generative, reasoning-based approaches, typically aligned via supervised fine-tuning (SFT) or preference optimization methods like Direct Preference Optimization (DPO). However, the increasing complexity of business rules and user queries exposes the inability of existing methods to endow models with robust reasoning capacity for long-tail and challenging cases. Efforts to address this via reinforcement learning strategies like Group Relative Policy Optimization (GRPO) often suffer from sparse terminal rewards, offering insufficient guidance for multi-step reasoning, which in turn slows convergence.
@@
TaoSR-AGRL was evaluated on large-scale real-world datasets and through online side-by-side human evaluations on Taobao Search. It consistently outperforms DPO and standard GRPO baselines in offline experiments, improving both relevance accuracy and rule adherence, while delivering more stable training. The model trained with TaoSR-AGRL has been successfully deployed in the main search scenario on Taobao, serving hundreds of millions of users.
@@ 1. Introduction
Large Language Models (LLMs)(yang2025qwen3technicalreport; deepseekai2025deepseekr1incentivizingreasoningcapability; grattafiori2024llama3herdmodels) have demonstrated remarkable reasoning capabilities in complex domains like math(havrilla2024teachinglargelanguagemodels; zhang2024llamaberrypairwiseoptimizationo1like; lai2024stepdpostepwisepreferenceoptimization) and code generation(hui2024qwen25codertechnicalreport; seed2025seedcoderletcodemodel). This has inspired their application to e-commerce search relevance, where methods like Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO)(rafailov2024directpreferenceoptimizationlanguage) are used to frame the task as a generative reasoning problem. However, these approaches often exhibit poor generalization on novel, reasoning-intensive long-tail samples, as they primarily learn to mimic patterns within the training distribution rather than acquiring robust, generalizable reasoning ability. To overcome these limitations in generalization and exploration, advanced reinforcement learning (RL) strategies like Group Relative Policy Optimization (GRPO)(shao2024deepseekmathpushinglimitsmathematical) are a promising direction. Yet, when applied to our context, their effectiveness is severely undermined by the critical challenge of reward sparsity(cui2025processreinforcementimplicitrewards; 10.5555/3600270.3600692; ecoffet2019go). This problem reflects both the intrinsic difficulty of the task and the suboptimal policy initialization: complex long-tail queries demand long, precise chains of reasoning under multiple constraints, making rewarding trajectories inherently rare and difficult to discover; meanwhile, the over-specialization induced by SFT or DPO yields a rigid, narrow exploratory prior, drastically reducing the already low probability of reaching valid reasoning paths. The confluence of these factors leads to exploration stagnation, ultimately crippling both training efficiency and model robustness on challenging unseen instances.
@@
To address the aforementioned issues, we propose an Adaptive Guided Reinforcement Learning Framework in Taobao Search Relevance (TaoSR-AGRL), a reinforcement learning framework for e-commerce relevance search. Its core idea is to alleviate the exploration bottleneck on hard, long-tail samples and enhance the model’s reasoning abilities through rule-aware, fine-grained reward shaping and adaptive guided trajectory resampling. TaoSR-AGRL introduces two key modules: (1) Rule-aware Reward Shaping: This module maps the complex rule-following process to a decomposed, weighted, multi-dimensional reward signal. It focuses on key dimensions of relevance reasoning, such as label, reasoning, and format. This transforms the originally sparse, all-or-nothing objective into a dense signal that encourages faithful reasoning. (2) Adaptive Guided Replay: For samples where the initial rollout fails to generate even partially correct reasoning trajectories, we design an adaptive replay mechanism. This mechanism identifies the model’s weaknesses revealed by Rule-aware Reward Shaping and uses the corresponding ground-truth labels as guidance to perform a second rollout, helping the model explore higher-value reasoning paths.
@@
We conduct our research on top of our previous work TaoSR1(dong2025taosr1thinkingmodelecommerce), using a proprietary relevance-focused large model, Tbstar, as the base model. It employs a Mixture-of-Experts (MoE) architecture with 42 billion total parameters and 3.5 billion activated parameters. We construct a test set containing four types of challenging queries to systematically evaluate TaoSR-AGRL. Offline experiments demonstrate that TaoSR-AGRL significantly outperforms baseline methods such as DPO, and GRPO. Furthermore, online side-by-side human evaluations show that the framework achieves consistent improvements in real-world business scenarios. Experiments and ablation studies confirm the effectiveness of Rule-aware Reward Shaping and Adaptive Guided Replay in mitigating reward sparsity on long-tail samples and enhancing the model’s reasoning capabilities. The resulting model has been deployed on Taobao, delivering significant business benefits. The main contributions of this paper are as follows:
@@ 2.1. Search Relevance
The modeling of search relevance has evolved from classical statistical models like TF-IDF (AIZAWA200345) and BM25 (10.1145/1645953.1646237), which rely on lexical matching, to the Learning to Rank (LTR) paradigm. LTR introduced models such as RankNet (10.1145/1102351.1102363) and LambdaMART (Burges2010FromRT), shifting from handcrafted functions to machine-learned optimization of IR metrics. The deep learning era emphasized semantic understanding over lexical overlap. After initial efforts with shallow networks like DSSM (huang2013learning), the Transformer architecture (vaswani2023attentionneed) and pre-trained models like BERT (devlin-etal-2019-bert) established two dominant architectural patterns: the efficient bi-encoder for scalable retrieval (karpukhin-etal-2020-dense) and the high-precision cross-encoder for re-ranking (nogueira2020passagererankingbert), presenting distinct approaches to balance computational efficiency with model effectiveness. More recently, Large Language Models (LLMs) have initiated a paradigm shift, proposing a complementary perspective that frames relevance not merely as a matching problem, but as a reasoning task. Frameworks like LREF (Tang_2025) and TaoSR1 (dong2025taosr1thinkingmodelecommerce) prompt LLMs to reason about e-commerce relevance, leveraging modern alignment techniques such as SFT, DPO (rafailov2024directpreferenceoptimizationlanguage), and GRPO (shao2024deepseekmathpushinglimitsmathematical). Despite their power, a critical limitation remains: reward sparsity (Zhang2025GRPOLEADAD). This issue is particularly acute in e-commerce search, as extensive alignment (SFT, DPO) for this singular task can over-specialize the model, constraining the exploratory capacity needed for GRPO to navigate the long-tail distribution. Our work, TaoSR-AGRL, introduces a novel optimization paradigm to specifically address this gap.
@@ 2.2. RLVR for LLM Reasoning
The paradigm for enhancing complex reasoning in Large Language Models (LLMs) has transitioned from Supervised Fine-Tuning (SFT) to Reinforcement Learning (RL). SFT excels at mimicking demonstrations but suffers from exposure bias and limited generalization (NIPS2015_e995f98d; chu2025sftmemorizesrlgeneralizes). In contrast, RL fosters superior generalization and exploration through direct environmental feedback (zhang2025onpolicyrlmeetsoffpolicy). A key development is Reinforcement Learning from Verifiable Rewards (RLVR), which underpins frontier models like OpenAI-o1 (openai2024openaio1card) and DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability). Algorithms such as GRPO use simple, outcome-based rewards to improve reasoning (zeng2025simplerlzooinvestigatingtamingzero). However, the efficacy of such methods is substantially limited in long-horizon tasks due to reward sparsity (nan2025ngrponegativeenhancedgrouprelative; quadros2025llmdrivenintrinsicmotivationsparse; tran2025exploitingtreestructurecredit). The low probability of discovering a correct solution through random exploration in such problems renders credit assignment highly challenging and hinders learning efficiency.
To address reward sparsity, prior work has explored two main avenues. The first, process supervision, densifies rewards by crediting intermediate steps. This is achieved by training Process Reward Models (PRMs) (nye2021workscratchpadsintermediatecomputation; wang2023selfconsistencyimproveschainthought; li2025enhancingreasoningprocesssupervision) or using an “LLM-as-a-Judge” (Hosseini2024VSTaRTV; Saha2025LearningTP), though the latter can face robustness and consistency issues (Gu2024ASO). The second avenue, external guidance, simplifies the learning task via hints from a superior LLM (nath2025adaptiveguidanceacceleratesreinforcement; zhang2025critiquegrpoadvancingllmreasoning) or partial reasoning trajectories (zhang2025stephintmultilevelstepwisehints). Both strategies, however, incur significant computational overhead or rely on external “oracle” models. In contrast, our work differs by introducing a lightweight, on-demand guidance mechanism. It provides targeted assistance only when needed, eliminating external dependencies and fostering autonomous, efficient learning with minimal computational cost.
@@ 3.1. Data Construction
We first identify four challenging query categories on which existing models perform poorly: negation, affordable alternatives, question-answering (Q&A), and knowledge-based queries. These initially filtered query logs were then submitted to a professional annotation team for fine-grained annotation based on the query-item relevance. The output of this process is formalized as tuples of <query, item, label_category, label_attribution, label_relevance>.
@@
To prioritize hard cases that provide the most informative training signal, we adopt a Difficulty-Aware Dynamic Sampling strategy, following the general principle in (yu2025dapoopensourcellmreinforcement). The base model performs multiple rounds of offline prediction over the training dataset, generating diverse predicted trajectories for each instance. Samples that are consistently predicted entirely incorrectly are then subjected to manual inspection, which reveals that their annotation accuracy is considerably low, indicating a substantial level of label noise in this subset. Therefore, our sampling process is designed to remove samples predicted either perfectly correctly or consistently completely incorrectly, as these extreme cases tend to be dominated by label noise or trivial patterns. The retained subset thus consists of medium-difficulty instances, which strike a balance between informativeness and annotation reliability, and are most effective for improving the model’s reasoning capacity.
@@ 3.2. Rule-aware Reward Shaping
The fidelity of the intermediate CoT is paramount for robust reasoning. However, a critical challenge arises because the aggregation mechanism of the derivation rules can mask underlying errors in the reasoning path. For example, consider a case where the correct reasoning path is “Category: Related” and “Attribute: Irrelevant”. According to the rules, this combination correctly yields a final relevance of 2-Mismatch. A flawed model, however, might correctly identify the category as “Related” but misjudge the attribute as “Mismatch”. Crucially, despite the incorrect intermediate step, the derivation rule would still lead the model to the same, correct final answer. When the model is optimized solely on the reward from this final outcome, it is incentivized to find such shortcuts rather than master the correct reasoning process. This is a classic case of reward hacking(skalse2025definingcharacterizingrewardhacking), which leads to two severe consequences: (1) Impaired Generalization: The model fails to learn the genuine reasoning logic, making it brittle when encountering out-of-distribution query-item pairs; and (2) Reduced Trustworthiness and Interpretability: A low-quality reasoning chain hinders error attribution and decision traceability in practical applications.
@@
To address these issues, we designed Rule-aware Reward Shaping. Its core idea is to decompose the single, sparse final reward signal into a dense, composite reward that provides fine-grained, multi-dimensional supervision across the entire reasoning chain. This mechanism is designed to ensure that the model not only produces the correct output but also adheres to a correct and interpretable reasoning path. Specifically, we deconstruct the total reward signal into three orthogonal components.
@@
  • Label Reward: This reward enforces the factual faithfulness of the reasoning chain by anchoring its key deductions to ground truth. It provides distinct signals for the correctness of the predicted category, attribute, and final relevance labels, ensuring the model’s conclusions are factually sound.
@@
  • Reasoning Reward: This reward enforces the logical faithfulness of the reasoning process itself, independent of the labels’ correctness. It is composed of two critical components: (1) Rule Adherence, which penalizes any step-wise deduction that violates the predefined derivation rules, and (2) Self-Consistency, which verifies that the model’s derived conclusion matches its initial response, preventing self-contradiction.
@@
  • Format Reward: This reward enforces the structural faithfulness of the model’s output, ensuring its syntactic integrity. It operates as a binary signal, providing a positive reward only if the generated CoT strictly conforms to the required structured format, thereby penalizing any malformed generation.
@@
To prevent the model from earning unearned rewards on flawed or invalid trajectories, we introduce a validity gating mechanism. This gate acts as a prerequisite for all fine-grained rewards, ensuring that credit is only assigned to reasoning paths that meet two fundamental criteria: structural integrity (correct format) and conclusive correctness (correct final relevance). This design establishes a natural curriculum learning(soviany2022curriculum). The model is first compelled to master the primary task of generating well-formed and correct final answers. Only after achieving this baseline competence does it receive the detailed, step-by-step guidance necessary to refine the faithfulness of its internal reasoning process. This staged approach efficiently channels the model’s learning, preventing it from getting lost optimizing flawed logic. This principle is formalized in our final reward function $R$:

### v3

@@ Abstract.
Query-product relevance prediction is fundamental to e-commerce search and has become even more critical in the era of AI-powered shopping, where semantic understanding and complex reasoning directly shape the user experience and exert an indirect, yet substantial impact on business conversion. Large Language Models (LLMs) enable generative, reasoning-based approaches, typically aligned via supervised fine-tuning (SFT) or preference optimization methods like Direct Preference Optimization (DPO). However, the increasing complexity of business rules and user queries exposes the inability of existing methods to endow models with robust reasoning capacity for long-tail and challenging cases. Efforts to address this via reinforcement learning strategies like Group Relative Policy Optimization (GRPO) often suffer from sparse terminal rewards, offering insufficient guidance for multi-step reasoning, which in turn slows convergence.
@@
TaoSR-AGRL was evaluated on large-scale datasets and through online evaluations on Taobao Search. It consistently outperforms DPO and GRPO baselines, improving relevance accuracy and rule adherence, with measurable gains in user engagement and stable training dynamics. The model has been deployed on Taobao, serving hundreds of millions of users.
@@ 1. Introduction
Large Language Models (LLMs)(Yang et al., 2025; DeepSeek-AI et al., 2025; Grattafiori et al., 2024) have demonstrated remarkable reasoning capabilities in complex domains like math(Havrilla et al., 2024; Zhang et al., 2024; Lai et al., 2024) and code generation(Hui et al., 2024; Seed et al., 2025). This has inspired their application to e-commerce search relevance, where methods like Supervised Fine-tuning (SFT) and Direct Preference Optimization (DPO)(Rafailov et al., 2024) are used to frame the task as a generative reasoning problem. However, these approaches often exhibit poor generalization on novel, reasoning-intensive long-tail samples, as they primarily learn to mimic patterns within the training distribution rather than acquiring robust, generalizable reasoning ability. To overcome these limitations in generalization and exploration, advanced reinforcement learning (RL) strategies like Group Relative Policy Optimization (GRPO)(Shao et al., 2024) represent a promising direction. Yet, when applied to our context, their effectiveness is severely undermined by the critical challenge of reward sparsity(Cui et al., 2025; Devidze et al., 2022; Ecoffet et al., 2019). This problem reflects both the intrinsic difficulty of the task and the suboptimal policy initialization: complex long-tail queries demand long, precise chains of reasoning under multiple constraints, making rewarding trajectories inherently rare and difficult to discover; meanwhile, the over-specialization induced by SFT or DPO yields a rigid, narrow exploratory prior, drastically reducing the already low probability of reaching valid reasoning paths. The confluence of these factors leads to exploration stagnation, ultimately crippling both training efficiency and model robustness on challenging unseen instances.
@@
To address the aforementioned issues, we propose an Adaptive Guided Reinforcement Learning Framework in Taobao Search Relevance (TaoSR-AGRL), a reinforcement learning framework for e-commerce relevance search. Its core idea is to alleviate the exploration bottleneck on hard, long-tail samples and enhance the model’s reasoning abilities through rule-aware, fine-grained reward shaping and adaptive guided trajectory resampling. TaoSR-AGRL introduces two key modules. First, Rule-aware Reward Shaping transforms the complex rule-following process into a decomposed, weighted, multi-dimensional reward signal. By explicitly modeling dimensions such as label accuracy, reasoning logic, and formatting constraints, it converts the sparse, binary objective into a dense signal that guides faithful reasoning. Second, Adaptive Guided Replay implements a corrective replay mechanism for samples where the initial rollout fails to produce valid reasoning trajectories. This mechanism identifies the model’s weaknesses revealed by Rule-aware Reward Shaping and uses the corresponding ground-truth labels as guidance to perform a second rollout, helping the model explore higher-value reasoning trajectories.
@@
We conduct our research on top of our previous work TaoSR1(Dong et al., 2025), using a proprietary relevance-focused large model, TbStar, as the base model. It employs a Mixture-of-Experts (MoE) architecture with 42 billion total parameters and 3.5 billion activated parameters. We construct a test set containing four types of challenging queries to systematically evaluate TaoSR-AGRL. Offline experiments demonstrate that TaoSR-AGRL significantly outperforms baseline methods such as DPO, and GRPO. Furthermore, both online side-by-side human evaluations and A/B testing demonstrate that the framework achieves consistent improvements in retrieval relevance and user engagement in real-world business scenarios. Experiments and ablation studies confirm the effectiveness of Rule-aware Reward Shaping and Adaptive Guided Replay in mitigating reward sparsity on long-tail samples and enhancing the model’s reasoning capabilities. The resulting model has been deployed on Taobao, delivering significant business benefits. The main contributions of this paper are as follows:
@@ 2.1. Search Relevance
The modeling of search relevance has evolved from classical statistical models like TF-IDF (Aizawa, 2003) and BM25 (Svore and Burges, 2009), which rely on lexical matching, to the Learning to Rank (LTR) paradigm. LTR introduced models such as RankNet (Burges et al., 2005) and LambdaMART (Burges, 2010), shifting from handcrafted functions to machine-learned optimization of IR metrics. The deep learning era emphasized semantic understanding over lexical overlap. After initial efforts with shallow networks like DSSM (Huang et al., 2013), the Transformer architecture (Vaswani et al., 2023) and pre-trained models like BERT (Devlin et al., 2019) established two dominant architectural patterns: the efficient bi-encoder for scalable retrieval (Karpukhin et al., 2020) and the high-precision cross-encoder for re-ranking (Nogueira and Cho, 2020), presenting distinct approaches to balance computational efficiency with model effectiveness. More recently, Large Language Models (LLMs) have initiated a paradigm shift, proposing a complementary perspective that frames relevance not merely as a matching problem, but as a reasoning task. Frameworks like LREF (Tang et al., 2025) and TaoSR1 (Dong et al., 2025) prompt LLMs to reason about e-commerce relevance, leveraging modern alignment techniques such as SFT, DPO (Rafailov et al., 2024), and GRPO (Shao et al., 2024). Despite their power, a critical limitation remains: reward sparsity (Zhang and Zuo, 2025). This issue is particularly acute in e-commerce search, as extensive alignment (SFT, DPO) for this singular task can over-specialize the model, constraining the exploratory capacity needed for GRPO to navigate the long-tail distribution. Our work, TaoSR-AGRL, introduces a novel optimization paradigm to specifically address this gap.
@@ 2.2. RLVR for LLM Reasoning
The paradigm for enhancing complex reasoning in Large Language Models (LLMs) has transitioned from Supervised Fine-Tuning (SFT) to Reinforcement Learning (RL). SFT excels at mimicking demonstrations but suffers from exposure bias and limited generalization (Bengio et al., 2015; Chu et al., 2025). In contrast, RL fosters superior generalization and exploration through direct environmental feedback (Zhang et al., 2025b). A key development is Reinforcement Learning from Verifiable Rewards (RLVR), which underpins frontier models like OpenAI-o1 (OpenAI et al., 2024) and DeepSeek-R1 (DeepSeek-AI et al., 2025). Algorithms such as GRPO use simple, outcome-based rewards to improve reasoning (Zeng et al., 2025). However, the efficacy of such methods is substantially limited in long-horizon tasks due to reward sparsity (Nan et al., 2025; Quadros et al., 2025; Tran et al., 2025). The low probability of discovering a correct solution through random exploration in such problems renders credit assignment highly challenging and hinders learning efficiency.
To address reward sparsity, prior work has explored two main avenues. The first, process supervision, densifies rewards by crediting intermediate steps. This is achieved by training Process Reward Models (PRMs) (Nye et al., 2021; Wang et al., 2023; Li et al., 2025) or using an “LLM-as-a-Judge” (Hosseini et al., 2024; Saha et al., 2025), though the latter can face robustness and consistency issues (Gu et al., 2024). The second avenue, external guidance, simplifies the learning task via hints from a superior LLM (Nath et al., 2025; Zhang et al., 2025c) or partial reasoning trajectories (Zhang et al., 2025a). Both strategies, however, incur significant computational overhead or rely on external “oracle” models. In contrast, our work differs by introducing a lightweight, on-demand guidance mechanism. It provides targeted assistance only when needed, eliminating external dependencies and fostering autonomous, efficient learning with minimal computational cost.
@@ 3.1. Data Construction
We first identify four challenging query categories on which existing models perform poorly: negation, affordable alternatives, question-answering (Q&A), and knowledge-based queries. These initially filtered query logs were then submitted to a professional annotation team for fine-grained annotation based on the query-item relevance. The output of this process is formalized as tuples of <query, item, label_category, label_attribute, label_relevance>.
@@
To prioritize hard cases that provide the most informative training signal, we adopt a Difficulty-Aware Dynamic Sampling strategy, following the general principle in (Yu et al., 2025). The base model performs multiple rounds of offline prediction over the training dataset, generating diverse predicted trajectories for each instance. Samples that are consistently predicted entirely incorrectly are then subjected to manual inspection, which reveals that their annotation accuracy is considerably low, indicating a substantial level of label noise in this subset. Therefore, our sampling process is designed to remove samples predicted either perfectly correctly or consistently completely incorrectly, as these extreme cases tend to be dominated by label noise or trivial patterns. The retained subset thus consists of medium-difficulty instances, which strike a balance between informativeness and annotation reliability, and are most effective for improving the model’s reasoning capacity.
@@ 3.2. Rule-aware Reward Shaping
The fidelity of the intermediate Chain-of-Thought (CoT) is paramount for robust reasoning. However, a critical challenge arises as the aggregation mechanism of derivation rules often masks underlying errors. Consider an instance where the ground truth reasoning involves identifying the category as Related and the attribute as Irrelevant. Under standard protocols, this combination maps to a final 2-Mismatch label. A flawed model, however, might correctly classify the category as Related while erroneously labeling the attribute as Mismatch. Crucially, despite this intermediate hallucination, the derivation logic still produces the correct final label. When the model is optimized solely on the reward from this final outcome, it is incentivized to find such shortcuts rather than master the correct reasoning process. This is a classic case of reward hacking(Skalse et al., 2025), yielding two severe consequences. First, it causes impaired generalization because the model fails to learn the genuine reasoning logic, rendering it brittle against out-of-distribution query-item pairs. Second, it leads to reduced trustworthiness and interpretability, as low-quality reasoning chains hinder error attribution and decision traceability in practical applications.
@@
To address these issues, we propose Rule-aware Reward Shaping. Its core idea is to decompose the single, sparse final reward signal into a dense, composite reward that provides fine-grained, multi-dimensional supervision across the entire reasoning chain. More generally, this realizes a transferable principle for rule-driven hierarchical decision tasks, where a final decision is composed from intermediate sub-decisions via explicit rules. In such settings, shaping rewards at each level enables precise credit assignment and discourages shortcut behaviors that exploit rule aggregation. This design ensures that the model not only produces the correct output but also follows a faithful and interpretable reasoning path. Specifically, we formulate the total reward via three orthogonal components:
@@
  • Label Reward: This component enforces the factual faithfulness of the reasoning chain by anchoring deductions to ground truth. It provides fine-grained supervision for the alignment of predicted category, attribute, and relevance labels, ensuring the final conclusion is factually sound.
@@
  • Reasoning Reward: This component enforces the logical faithfulness of the reasoning process, independent of label correctness. It comprises two sub-mechanisms: (1) Rule Adherence, which penalizes step-wise deductions that violate predefined derivation rules; and (2) Self-Consistency, which verifies that the final derived conclusion aligns with the initially generated prediction.
@@
  • Format Reward: This component enforces the structural faithfulness of the output, preserving syntactic integrity. It functions as a binary gating signal, yielding a positive reward solely when the generated CoT strictly adheres to the required schema, thereby filtering out malformed sequences.
@@
To prevent the model from earning unearned rewards on flawed or invalid trajectories, we introduce a validity gating mechanism. This gate acts as a prerequisite for all fine-grained rewards, ensuring that credit is only assigned to reasoning paths that meet two fundamental criteria: structural integrity (correct format) and conclusive correctness (correct final relevance). This design establishes a natural curriculum learning(Soviany et al., 2022). The model is first compelled to master the primary task of generating well-formed and correct final answers. Only after achieving this baseline competence does it receive the detailed, step-by-step guidance necessary to refine the faithfulness of its internal reasoning process. This staged approach efficiently channels the model’s learning, preventing it from getting lost optimizing flawed logic. This principle is formalized in our final reward function $R$: